OpenStudy (anonymous):

In the problem 15x^3y^5/ 20x^7y Chloe got 3x^4y^5/4 What are her 2 mistakes and what should the answer be? someone please help me !!!!!!!!!

OpenStudy (nurali):

\[\frac{ 15x^3y^5 }{ 20x^7y }\] \[\frac{ 3x ^{3-7}y ^{5-1} }{ 4 }\] \[\frac{ 3x ^{-4}y^4 }{ 4 }\] \[\frac{ 3y^4 }{ 4x^4 }\]